OpenStudy (anonymous):

A playground is being designed where children can interact with their friends in certain combination i have the answer i need to know how to find the recursive equation that represents the pattern

OpenStudy (anonymous):

an = an – 1 + 3(n – 1) -------------------------------- verify: a3=9 a4 = a3+3(4-1) a4 = 9 +3(3) = 18

OpenStudy (anonymous):

18=answer

OpenStudy (anonymous):

Post the entire question.

OpenStudy (anonymous):

A playground is being designed where children can interact with their friends in certain combinations. If there is 1 child, there can be 0 interactions. If there are 2 children, there can be only 1 interaction. If there are 3 children, there can be 4 interactions. If there are 4 children, there can be 9 interactions. Which recursive equation represents the pattern?

OpenStudy (anonymous):

Answer?